# Thin films spreading under the action of surface acoustic waves

### v1

@@ 3 Model
with $\phi_{s}$ given by Eq. (15) and $\phi_{g}(z)=\rho_{0}gz$, corresponding to a gravitational body force, $\mathbf{F}_{g}=-\nabla\phi_{g}=-\rho_{0}g\,\hat{\mathbf{k}}$, where $\hat{\mathbf{k}}$ is a unit vector in the out-of-plane direction.
In this work, we consider incompressible, low Reynolds number free surface flows under the long-wave approximation so that all inertial terms on the left-hand side of Eqs. (16) are negligible with respect to the viscous and pressure terms. Carrying out a standard long-wave expansion of (16) and imposing the usual stress boundary conditions at the free surface (vanishing tangential stress and normal stress balanced by surface tension, see e.g. Kondic (2003) for details), we obtain the following fourth-order differential equation for the film height, $z=h(x,y,t)$,
@@
We nondimensionalize the problem using an arbitrary length scale, $\ell$, a timescale $t_{c}$ such that leading order terms balance in (18), and a pressure scale $p_{c}$ based on the capillary contribution in (19), so that we have
@@
${\rm Bo}=\frac{\rho_{0}\ell^{2}}{\gamma}g=\frac{\ell^{2}}{a^{2}},\qquad{\cal S}=\frac{\ell}{\gamma}\phi_{0},$ (24)
where $a=\sqrt{\gamma/(\rho_{0}g)}$ is the capillary length and $\phi_{0}$ was defined in (15). In our simulations, motivated by a typical drop size in the experiments, we set $\ell=1$ mm. For simplicity, we omit the tilde from now on so that all variables will be dimensionless unless specific units are explicitly written, mainly when comparison with experimental data is required.
4 Theoretical results
In this section, we discuss the predictions of the model developed in Sec. 3, and compare them with the experimental results. In Sec. 4.1 we present results for a chosen reference case, and then in Sec. 4.2, we discuss the influence of model parameters, in particular the acoustic amplitude, on the results. In Sec. 4.3 we discuss (approximate) traveling-wave solutions. All simulations are carried out in two physical dimensions, assuming translational invariance in the transverse $y$-direction. The discussion of the influence of the three-dimensional configuration will be presented elsewhere.
4.1 Main features of the results
@@
$h(x,0)=\left\{\begin{array}[]{ll}h_{p},&x<x_{r}(0),\\ \left(h_{d}-h_{p}\right)\left[1-\left(\frac{x-x_{d}}{r_{d}}\right)^{2}\right]+h_{p},&x_{r}(0)\leq x\leq x_{f}(0),\\ h_{p},&x>x_{f}(0),\end{array}\right.$ (26)
where $x_{r}(0)=x_{d}-r_{d}$ and $x_{f}(0)=x_{d}+r_{d}$ are the initial positions of the rear and front contact lines of the drop, respectively (see Fig. 5). Here, $h_{p}$ ($\ll h_{d}$) is the thickness of a precursor film, introduced for numerical reasons in order to avoid the stress singularity associated with a moving contact line. It is well known (see, e.g. Diez & Kondic (2002)), that the step size $\Delta x$ in the spatial discretization should be similar to $h_{p}$ to ensure numerical convergence, and therefore we use $\Delta x=h_{p}$ in our simulations. The specific value assigned to $h_{p}$ has only a minor influence on the results; e.g., if $h_{p}$ is doubled or halved, the change of spreading speed or drop height is just 1-2%. All reported results are obtained using $h_{p}=0.01$.
Recalling the introduction of the reference $x$-coordinate origin, $x_{\rm ref}(t)$, in Sec. 3, we note that the SAW experiences minimal attenuation in regions of the actuator where fluid is not present, with significant attenuation only under the bulk droplet. To model this, we assume that the acoustic potential, $\phi_{s}$, is constant (no attenuation) in regions where the film height is smaller than some critical thickness $h^{*}$, and attenuates where $h>h^{*}$,
@@
where $x_{1}^{\ast}$, $x_{2}^{\ast}$ are defined by $h(x_{1}^{\ast},t)=h(x_{2}^{\ast},t)=h^{\ast}$ (this form of $\phi_{s}$ corresponds to the choice $x_{\rm ref}(t)=x_{1}^{*}(t)$). This model is further motivated by the known fact that the Eckart streaming—the acoustic streaming that pushes the oil film along the path of SAW in this work—governs the flow when the oil film is sufficiently thick. Close to the solid surface the flow is governed by a boundary layer flow (known as the Schlichting streaming to some and Rayleigh streaming to others), generated by the SAW in the solid. These mechanisms are observed to support the spreading of $\lambda_{\rm oil}/4$ thick oil films opposite the path of the SAW. However, thicker oil films are governed by Eckart streaming and spread along the path of the SAW. Hence, we account for Eckart streaming-induced acoustic forcing where the film thickness is greater than $h^{\ast}=\lambda_{\rm oil}/(4\ell)$  (Rezk et al., 2012, 2014; Altshuler & Manor, 2015, 2016).
[…]
@@
[…]
Figure 6(a) shows the evolution of the thickness profile obtained by solving Eq. (21) for the parameter values in Table 1, subject to the initial condition given by Eq. (26), and boundary conditions fixing the film thickness to $h_{p}$ and requiring vanishing derivatives at the domain boundaries. We observe that the simulated drop settles to a steady shape, which translates uniformly. The steady shape of the bulk droplet has a pronounced “foot” at its rear; though we do not yet fully understand the details of why this forms, it is consistent with the overall drop asymmetry that develops, seen in Fig. 2. We discuss these features of the results further below.
@@
Figure 6: (a) Thickness profiles from $t=0$ (red line) to $t=6$ s every $\Delta t=0.5$ s for SAW amplitude $A=1.30\,$nm. Here, we use $h^{\ast}=17\mu{\rm m}/h_{c}\approx(\lambda_{\rm oil}/4)/h_{c}$. (b) Comparison between simulated (solid blue line) and experimental (black symbols) thickness profiles at $t=2$ s.
4.2 Parametric dependence
@@
In this section, we separate the discussion into two parts: first, we consider early-time behavior, and then long-time evolution. Figure 7 shows the computational results for the maximum drop height $h_{\rm max}$ for different values of the acoustic amplitude, $A$. We show the results for two different initial conditions, with the purpose of illustrating the relevance of the initial drop shape. Focusing first on the initial condition used for Fig. 6 (solid curves in Fig. 7) with initial drop height of $h_{d}\approx 0.5$ mm obtained from the experiments (see also Eq. (25)), we note different behavior for small and large values of $A$, as could be anticipated from the experimental results of Fig. 4(b). For the larger $A$-values, $h_{\rm max}$ initially increases rapidly but then decreases. For the two smallest $A$-values, this initial rapid increase is not observed. In all cases, for longer times, $h_{\rm max}$ slowly decreases towards a constant ($A$-dependent) value, although it does not completely saturate for the times shown. One interpretation of these results is that for large $A$, the acoustic forcing dominates capillary stresses and gravity, leading to an increase of the drop height, while for smaller $A$, the acoustic forcing is not strong enough to lead to initial drop height increase. For the other initial condition (smaller initial drop height, shown by dashed lines in Fig. 7), the details are different, but we again observe an initial increase of $h_{\rm max}$ for sufficiently large values of $A$, followed by (in this case) a slight decrease towards constant $A$-dependent values at large times.
@@
Figure 7: Maximum fluid thickness, $h_{\rm max}$, as a function of time at different SAW amplitudes, $A$, for two initial conditions of the drop: $r_{d}=3.2$ mm (solid lines) and $r_{d}=3.9$ mm (dashed lines). The cross–sectional fluid area, $A_{d}$, is the same, but the latter has a lower initial maximum thickness than the former (see Eq. (25)). The values of are $A$ are: $2.00$ nm (black), $1.69$ nm (blue), $1.30$ nm (red), $1.04$ nm (magenta), $0.65$ nm (brown), $0.20$ nm (green) (same colors correspond to same values of $A$).
@@
Next, we shift our attention to long-time behavior, focusing first on the dependence of drop profiles on SAW amplitude $A$. Figure 8 shows the long-time drop profiles together with the pressure specified by Eq. (22) for four values of $A$. While the maximum drop height increases with $A$, as expected, we note that the thickness of the ‘foot’ directly behind the front increases with $A$ as well (varying between approximately 50 and 100 $\mu$m for the considered values of $A$). In addition, we note a very thin (on the order of the precursor scale, but a little larger) trailing film left behind the translating drop. While both the ‘foot’ and the thin trailing film deposited behind the drop still need to be fully analyzed and understood, we suspect that these features of the results are related to the pressure (plotted by the red lines with values shown on the right-hand axis in Fig. 8); note that the ‘foot’ forms in the region where the pressure is O(1), shown by the dashed lines. On a different note, we point out the strong decrease of the pressure due to the SAW attenuation, as the film height increases from the precursor scale, as well as the pressure dip at the front contact line, caused by the rapidly changing curvature there.
@@
Figure 8: Simulated drop profiles (blue curves, left axis) and dimensionless pressure $\tilde{P}=P/p_{c}$ (red curves, right axis) at $t=6$ s for four different values of SAW amplitude $A$. The horizontal dashed line is at $\tilde{P}=1$.
Focusing next on the spreading speed, Fig. 9 illustrates that for long times both experiments and simulations show approximately constant spreading speed (drop front position $x_{f}$ increasing linearly in time), which seems to be a characteristic feature of the considered acoustic driving (note that we shift the experimental results in the figure to illustrate better the long time behavior; initial transients may account for differences at early times due to the particulars of the initial conditions). These results suggest that the spreading drop essentially translates without much of shape change for late times. Figure 10(a) confirms this expectation: this figure shows that the width of the spreading drop, ${\rm w}=x_{2}^{*}-x_{1}^{*}$, reaches a plateau and remains constant in the case of larger acoustic amplitudes $A$; on sufficiently long time scales (not shown in the figure), the $\rm w$-values obtained for smaller $A$-values reach a plateau as well.
Figure 9: Comparison between numerical (solid lines) and experimental (symbols) results for the SAW amplitude $A=1.69\,$nm (blue), 1.301.301.30\,1.30nm (red) and 1.041.041.04\,1.04nm (magenta). We have added arbitrary $x$–shifts to the simulation results to facilitate the comparison of the slope of $x_{f}$ versus $t$ between experiments and simulations for long times.
Figure 10: (a) Numerically calculated drop width, $\rm w$, as a function of time for several values of $A$: $=0.20$ nm (green), $=0.65$ nm (brown), $=1.04$ nm (magenta), $=1.30$ nm (red), $=1.69$ nm (blue), and $=2.00$ nm (black). Note that the two smallest values of $A$ have not reached an asymptotic value of $\rm w$ for $t=6$ s, but will do for later times. (b) Long time values of the front speed, ${\rm v}_{f}$ (in mm/s, blue filled circles), maximum thickness, $h_{\rm max}$ (mm, red hollow squares), and drop width, $\rm w$ (cm, black filled squares), as functions of $A$. The points correspond to the raw data, and the solid lines connecting them guide the eye.
To summarize our numerical results, Fig. 10(b) plots the numerically-predicted long-time values (reached for $t\gtrsim 6$ s) of the front velocity, ${\rm v}_{f}$, the maximum drop height, $h_{\rm max}$, and the drop width, ${\rm w}=x_{2}^{\ast}-x_{1}^{\ast}$, as a function of $A$. We note that, while ${\rm v}_{f}$ strongly increases with $A$, the dependence of $h_{\rm max}$ and $\rm w$ on $A$ is much weaker. These results suggest that a good qualitative description of the problem might be obtained by considering a traveling wave type of solution. Such an approach is discussed in the following section.
4.3 Traveling wave solution

### v2

@@ III Model
with $\phi_{s}$ given by Eq. (15) and $\phi_{g}(z)=\rho_{\rm{oil}}gz$, corresponding to a gravitational body force, $\mathbf{F}_{g}=-\nabla\phi_{g}=-\rho_{\rm{oil}}g\,\hat{\mathbf{k}}$, where $\hat{\mathbf{k}}$ is a unit vector in the out-of-plane direction.
We now briefly discuss the assumptions of the model presented so far. Our focus is on modeling drops of millimetric thickness, as considered in our experiments. Therefore, we focus on the Eckart streaming that dominates at these length scales; that is, the variation of the acoustic field in the solid caused by attenuation of SAW in the solid substrate, as discussed in the Introduction and as considered thus far in our presentation. We recall that, very close to the solid surface, there is a boundary layer flow governed by Schlichting and Rayleigh streaming, in which Eckart streaming is negligible. These Schlichting and Rayleigh mechanisms are observed to support spreading on lengthscales $\lambda_{\rm oil}/4$ or less [44, 45, 46, 47], which is much smaller than the drop thickness relevant to our experiments. These effects are therefore not included in our formulation, except that in our implementation (described in Sec. IV), we assume that the SAW experiences no attenuation in regions where the film thickness is smaller than this value so that Eckart streaming does not act in such regions.
We also ignore contributions from acoustic radiation pressure. To justify this approximation, we note that the stress associated with acoustic radiation pressure due to a traveling acoustic wave is characterized by [35] $\left(1/2+B/(4C)\right)\rho_{\text{oil}}\left(\omega A\right)^{2}$, where $B$ and $C$ are the Fox and Wallace coefficients for the non-linear adiabatic connection between pressure and fluid density. For our experiments, $\left(1/2+B/(4C)\right)\approx 3$, given that $B/C\approx 10$ for organic liquids at room temperature [54]. On the other hand, the characteristic stress associated with Eckart–type acoustic streaming is approximately $\left(l_{i}f/c_{\rm s}\right)\rho_{\text{oil}}\left(\omega A\right)^{2}$, noting again that $l_{i}$ and $f/c_{\rm s}$ are the attenuation length and the wavelength of the SAW. The product $l_{i}f/c_{\rm s}=\left(\rho_{\rm s}/\rho_{\text{oil}}\right)\left(c_{\rm s}/c_{\text{oil}}\right)\approx 16$, using the experimental values of the parameters given after Eq. (10). Hence, as long as the liquid drop is sufficiently thick to support Eckart streaming (as is the case in our experiments, where the drops are of millimetric thickness), and while its surface is curved so that acoustic resonance effects that may result in high acoustic radiation pressure levels are avoided, the stresses that contribute to flow due to the Eckart streaming mechanism are an order of magnitude larger than contributions from acoustic radiation pressure.
In this work, we consider incompressible, low Reynolds number free surface flows under the long-wave approximation so that all inertial terms on the left-hand side of Eqs. (16) are negligible with respect to the viscous and pressure terms. Carrying out a standard long-wave expansion of Eqs. (16) and imposing the usual stress boundary conditions at the free surface (vanishing tangential stress and normal stress balanced by surface tension, see e.g. [55] for details), we obtain the following fourth-order differential equation for the film height, $z=h(x,y,t)$,
@@
We nondimensionalize the problem using an arbitrary length scale, $\ell$, a timescale $t_{c}$ such that leading order terms balance in Eq. (18), and a pressure scale $p_{c}$ based on the capillary contribution in Eq. (19), so that we have
@@
${\rm Bo}=\frac{\rho_{\rm{oil}}\ell^{2}}{\gamma}g=\frac{\ell^{2}}{a^{2}},\qquad{\cal S}=\frac{\ell}{\gamma}\phi_{0},$ (24)
where $a=\sqrt{\gamma/(\rho_{\rm{oil}}g)}$ is the capillary length and $\phi_{0}$ was defined in Eq. (15). In our simulations, motivated by a typical drop size in the experiments, we set $\ell=1$ mm. For simplicity, we omit the tilde from now; all variables are dimensionless unless specific units are explicitly given, mainly when comparison with experimental data is required.
IV Theoretical results
In this section, we discuss the predictions of the model developed in Sec. III, and compare them with the experimental results. In Sec. IV.1, we present results for a chosen reference case, and then in Sec. IV.2, we discuss the influence of model parameters, in particular the acoustic amplitude, on the results. In Sec. IV.3 we discuss (approximate) traveling-wave solutions. All simulations are carried out in two physical dimensions, assuming translational invariance in the transverse $y$-direction. The discussion of three-dimensional effects will be presented elsewhere.
IV.1 Main features of the results
@@
$h(x,0)=\left\{\begin{array}[]{ll}h_{p},&x<x_{\rm r}(0),\\ \left(h_{d}-h_{p}\right)\left[1-\left(\frac{x-x_{d}}{r_{d}}\right)^{2}\right]+h_{p},&x_{\rm r}(0)\leq x\leq x_{\rm f}(0),\\ h_{p},&x>x_{\rm f}(0),\end{array}\right.$ (26)
where $x_{\rm r}(0)=x_{d}-r_{d}$ and $x_{\rm f}(0)=x_{d}+r_{d}$ are the initial positions of the rear and front contact lines of the drop, respectively (see Fig. 5). Here, $h_{p}$ ($\ll h_{d}$) is the thickness of a precursor film, introduced for numerical reasons in order to avoid the stress singularity associated with a moving contact line. It is well known (see, e.g. [56]), that the step size $\Delta x$ in the spatial discretization should be similar to $h_{p}$ to ensure numerical convergence, and therefore we use $\Delta x=h_{p}$ in our simulations. The specific value assigned to $h_{p}$ has only a minor influence on the results; e.g., if $h_{p}$ is doubled or halved, the change of spreading speed or drop height is just 1-2%. All reported results are obtained using $h_{p}=0.01$.
Recalling the introduction of the reference $x$-coordinate origin, $x_{\rm ref}(t)$, in Sec. III, we note that the SAW experiences minimal attenuation in regions of the actuator where fluid is not present, with significant attenuation only under the bulk droplet. To model this, we assume that the acoustic potential, $\phi_{s}$, is constant (no attenuation) in regions where the film height is smaller than some critical thickness $h^{*}$, and attenuates where $h>h^{*}$,
@@
where $x_{1}^{\ast}$, $x_{2}^{\ast}$ are defined by $h(x_{1}^{\ast},t)=h(x_{2}^{\ast},t)=h^{\ast}$ (this form of $\phi_{s}$ corresponds to the choice $x_{\rm ref}(t)=x_{1}^{*}(t)$). In line with our discussion in Sec. III, in simulations we choose the cutoff film thickness $h^{*}$ based on the thickness below which the Eckart streaming ceases to dominate the Schlichting and Rayleigh streaming, $h^{\ast}\approx(\lambda_{\rm oil}/4)/\ell=17\mu{\rm m}/\ell$.
We commence our analysis by considering the experimental case where an oil drop of a volume of $8$ mm3 is under the excitation of the 20 MHz SAW that supports a normal surface displacement amplitude of $A=1.3$ nm before attenuation, shown in Fig. 4. Table 1 gives the corresponding dimensional and dimensionless parameters (see also Eqs. (9) and (10)).
@@
$\ell=1$ mm $h_{p}=0.010$
$h_{\ast}=0.017$
Table 1: Dimensional parameters of the experiments and the corresponding dimensionless constants used in the simulations for an $8$ mm3 volume drop.
Figure 6(a) shows the evolution of the thickness profile obtained by solving Eq. (21) for the parameter values in Table 1, subject to the initial condition given by Eq. (26), and boundary conditions fixing the film thickness to $h_{p}$ and requiring vanishing derivatives at the domain boundaries. We observe that the simulated drop settles to a steady shape, which translates uniformly. The steady shape of the bulk droplet has a pronounced “foot” at its rear; though we do not yet fully understand the details of why this forms, it is consistent with the overall drop asymmetry that develops, see Fig. 2. We discuss these features of the results further below.
@@
Figure 6: (a) Thickness profiles from $t=0$ (red line) to $t=6$ s every $\Delta t=0.5$ s for $A=1.30\,$nm. Here, we use $h^{\ast}=17\mu{\rm m}/\ell\approx(\lambda_{\rm oil}/4)/\ell$. (b) Comparison between simulated (solid blue line) and experimental (black symbols) thickness profiles at $t=2$ s.
IV.2 Parametric dependence
@@
In this section, we separate the discussion into two parts: first, we consider early-time behavior, and then long-time evolution. Figure 7 shows the computational results for the maximum drop height $h_{\rm max}$ for different values of $A$. We show the results for two different initial conditions, with the purpose of illustrating the relevance of the initial drop shape. Focusing first on the initial condition used for Fig. 6 (solid curves in Fig. 7) with initial drop height of $h_{d}\approx 0.5$ obtained from the experiments (see also Eq. (25)), we note different behavior for small and large values of $A$, as could be anticipated from the experimental results of Fig. 4(b). For the larger $A$-values, $h_{\rm max}$ initially increases rapidly but then decreases. For the two smallest $A$-values, this initial rapid increase is not observed. In all cases, for longer times, $h_{\rm max}$ slowly decreases towards a constant ($A$-dependent) value, although it does not completely saturate for the times shown. One interpretation of these results is that for large $A$, the acoustic forcing dominates capillary stresses and gravity, leading to an increase of the drop height, while for smaller $A$, the acoustic forcing is not strong enough to lead to initial drop height increase. For the other initial condition (smaller initial drop height, shown by dashed lines in Fig. 7), the details are different, but we again observe an initial increase of $h_{\rm max}$ for sufficiently large values of $A$, followed by (in this case) a slight decrease towards constant $A$-dependent values at large times.
@@
Figure 7: Maximum fluid thickness, $h_{\rm max}$, as a function of time for two initial conditions of the drop: $r_{d}=3.2$ (solid lines) and $r_{d}=3.9$ (dashed lines) (in units of $\ell$). The cross–sectional fluid area, $A_{d}$, is the same, but the latter has a lower initial maximum thickness than the former (see Eq. (25)). The values of $A$ are: $2.00$ nm (black), $1.69$ nm (blue), $1.30$ nm (red), $1.04$ nm (magenta), $0.65$ nm (brown), $0.20$ nm (green).
@@
Next, we shift our attention to long-time behavior, focusing first on the dependence of drop profiles on SAW amplitude. Figure 8 shows the long-time drop profiles together with the dimensionless pressure specified by Eq. (22) for four values of $A$. While the maximum drop height increases with $A$, as expected, we note that the thickness of the ‘foot’ directly behind the front increases with $A$ as well (varying between approximately 50 and 100 $\mu$m for the considered values of $A$). In addition, we note a very thin trailing film, just slightly thicker than the precursor left behind the translating drop. While both the ‘foot’ and the thin trailing film still need to be fully analyzed and understood, we suspect that these features of the results are related to the pressure (plotted by the red lines with values shown on the right-hand axis in Fig. 8); note that the ‘foot’ forms in the region where the dimensionless pressure is O(1), shown by the dashed lines. On a different note, we point out the strong decrease of the pressure due to the SAW attenuation, as the film height increases from the precursor scale, as well as the pressure dip at the front contact line, caused by the rapidly changing curvature there.
@@
Figure 8: Simulated drop profiles (blue curves, left axis) and dimensionless pressure $\tilde{P}$ (red curves, right axis) at $t=6$ s for four different values of $A$. The horizontal dashed line is at $\tilde{P}=1$.
Focusing next on the spreading speed, Fig. 9 illustrates that for long times both experiments and simulations show approximately constant spreading speed (drop front position $x_{\rm f}$ increasing linearly in time), which seems to be a characteristic feature of the considered acoustic driving (note that we shift the experimental results in the figure to illustrate better the long time behavior; initial transients may account for differences at early times due to the particulars of the initial conditions). These results suggest that the spreading drop essentially translates without much shape change at late times. Figure 10(a) confirms this expectation: this figure shows that the width of the spreading drop, ${\rm w}=x_{2}^{*}-x_{1}^{*}$, reaches a plateau and remains constant in the case of larger acoustic amplitudes $A$; on sufficiently long time scales (not shown in the figure), the $\rm w$-values obtained for smaller $A$-values reach a plateau as well.
Figure 9: Comparison between numerical (solid lines) and experimental (symbols) results for $A=1.69\,$nm (blue), 1.301.301.30\,1.30nm (red) and 1.041.041.04\,1.04nm (magenta). We have added arbitrary $x$–shifts to the simulation results to facilitate the comparison of the slope of $x_{\rm f}$ versus $t$ between experiments and simulations for long times.
Figure 10: (a) Numerically calculated drop width, $\rm w$, as a function of time for several values of $A$: $A=0.20$ nm (green), $A=0.65$ nm (brown), $A=1.04$ nm (magenta), $A=1.30$ nm (red), $A=1.69$ nm (blue), and $A=2.00$ nm (black). Note that the two smallest values of $A$ have not reached an asymptotic value of $\rm w$ for $t=6$ s, but will do for later times. (b) Long time values of the front speed, ${\rm v}_{\rm f}$ (in mm/s, blue filled circles), maximum thickness, $h_{\rm max}$ (mm, red hollow squares), and drop width, $\rm w$ (cm, black filled squares), as functions of $A$. The points correspond to the raw data, and the solid lines connecting them guide the eye.
To summarize our numerical results, Fig. 10(b) plots the numerically-predicted long-time values (reached for $t\gtrsim 6$ s) of the front velocity, ${\rm v}_{\rm f}$, the maximum drop height, $h_{\rm max}$, and the drop width, ${\rm w}=x_{2}^{\ast}-x_{1}^{\ast}$, as a function of $A$. We note that, while ${\rm v}_{\rm f}$ strongly increases with $A$, the dependence of $h_{\rm max}$ and $\rm w$ on $A$ is much weaker. These results suggest that a good qualitative description of the problem might be obtained by considering a traveling wave type of solution. Such an approach is discussed in the following section.
IV.3 Traveling wave solution